Your firm currently has 1,932,460 shares of stock outstanding that sell for 34.29. Additionally, you've issued 27,824 bonds, with a par value of S1,000. They currently sell for 110.4% of par.

What is the firm's equity weight?

(Enter your answer as a percentage to 1 decimal place. ex: 12.3)
Answer #1

Total value of:

Equity=(1,932,460*34.29)=$66264053.4

Bonds=27,824*(1000*110.4%)=$30717696

Total value=(66264053.4+30717696)=$96981749.4

Hence equity weight=(66264053.4/96981749.4)

=68.3%(Approx).